# Extremum Seeking Tracking for Derivative-free Distributed Optimization

### v4

@@ Abstract
In this paper, we deal with a network of agents that want to cooperatively minimize the sum of local cost functions depending on a common decision variable. We consider the challenging scenario in which objective functions are unknown and agents have only access to local measurements of their local functions. We propose a novel distributed algorithm that combines a recent gradient tracking policy with an extremum-seeking technique to estimate the global descent direction. The joint use of these two techniques results in a distributed optimization scheme that provides arbitrarily accurate solution estimates through the combination of Lyapunov and averaging analysis approaches with consensus theory. We perform numerical simulations in a personalized optimization framework to corroborate the theoretical results.
@@ I Introduction
We organize the literature review in two blocks: collaborative/distributed extremum seeking and distributed zeroth-order/derivative-free optimization. Early references on distributed extremum for the so-called consensus optimization framework are [8] for a discrete-time setting and [9] for a continuous-time one. A proportional-integral extremum-seeking design technique is proposed in [10], while, in [11], the gradient is approximated through a real-time protocol. In [12], authors propose the use of the sliding mode to generate the dither signal at the base of the extremum seeking. More recently, in [13], a distributed stochastic extremum seeking scheme is proposed for a source localization problem, while in [14] a distributed scheme approximating a Newton method is proposed. As for constraint-coupled distributed optimization, in [15], a distributed extremum seeking control, based on evolutionary game theory, is designed for real-time resource allocation. In [16], instead, a distributed, continuous-time scheme based on sign-based consensus is designed. In [17], a Lie bracket technique and extremum-seeking are used for problems with linear constraints. In [18], resource allocation problems are addressed by an extremum-seeking scheme.
@@
As for distributed zeroth-order/derivative-free optimization, authors in [19] develop a zeroth-order scheme based on a one-point estimator and a gradient tracking policy. The work [20] instead proposes a zeroth-order algorithm based on a two-point estimator with a distributed gradient descent strategy and another one based on an $n$-point estimator with a gradient tracking policy. Authors in [21] propose a continuous-time gradient-free approach emulating a distributed gradient algorithm for which optimal asymptotic convergence is guaranteed. In [22], a sampled version of [21] is proposed. In [23], a continuous-time distributed algorithm based on random gradient-free oracles is proposed for convex optimization problems. In [24], the randomized gradient-free oracles introduced in [23] are used to build a gradient-free distributed algorithm in directed networks. Randomized gradient-free algorithms are used also in [25], where sequential Gaussian smoothing is used for non-smooth distributed convex constrained optimization. In [26], gradient-free optimization is addressed with the additional constraint that each agent can only transmit quantized information. The work in [27] instead develops a “directed-distributed projected pseudo-gradient” descent method for directed graphs. Paper [28] combines the gradient-free strategy of [23] with a saddle-point algorithm. Authors in [29] address an online constrained optimization problem by relying on the Kiefer-Wolfowitz algorithm to approximate the gradients, and [30] combines the estimation of the gradient via a “simultaneous perturbation stochastic approximation” technique with the so-called matrix exponential learning optimization method. In [31], a randomized gradient-free method is combined with a state-of-the-art distributed gradient descent approach for directed networks. In [32], an overview of zeroth-order methods based on a Frank-Wolfe framework is provided. Authors in [33] propose a distributed random gradient-free protocol to solve constrained optimization problems by using projection techniques. All the cited papers but [30] prove stability of the proposed schemes by using arguments not based on Lyapunov theory. We also note that [30] proposes an algorithm neither based on gradient tracking nor using extremum-seeking.
@@
This paper proposes a novel algorithm to solve a distributed optimization problem in which network agents can only evaluate their local cost function at a given point, but not its gradient. The proposed solution consists of a distributed protocol in which the (unavailable) local gradients are approximated through an extremum seeking scheme. The approximations of the gradients are used to feed a suitable tracking mechanism which, in turn, allows for steering the local solution estimates along approximations of the global descent direction. The distributed algorithm uses a further consensus action on the local solution estimates. The convergence of our scheme is proved through Lyapunov and averaging theories for discrete-time systems. It is worth mentioning that our scheme, together with the ones in [34, 35], is the only distributed extremum-seeking scheme proposed in discrete-time with the following distinctive features. The work [34] (i) does not address a consensus optimization problem and (ii) relies on consensus dynamics estimating the global cost, while ours estimates the global gradient. Instead, in [35], (i) the addressed consensus optimization problems have scalar decision variables and (ii) an extremum-seeking technique is combined with a distributed gradient algorithm, i.e., without a tracking mechanism. As for the literature on distributed zeroth-order/derivative-free methods, the closest work is [19] in which extremum seeking is not used and the global gradient is approximated via a randomized one-point policy. Consistent with the comparison table provided in [19], we highlight that there are no other distributed algorithms in the literature using 1-point gradient estimators. Indeed, although the distributed scheme in [36] uses 1-point gradient estimators, we remark that it is tailored for partition-based optimization, i.e., a simplified setup in which each local function depends only on the neighbors’ decision variables, see [3]. Also, our Lyapunov-based tools combined with averaging theory for discrete-time systems represent a distinctive feature in the algorithm analysis.
@@ II Problem Formulation
Each cost $f_{i}$ is $\mathcal{C}^{3}$ and has $\overline{L}_{i}$-Lipschitz continuous gradients. We denote $\overline{L}=\max\{\overline{L}_{1},\dots,\overline{L}_{N}\}$. $\square$
Assumption 2 ensures the existence of a unique solution $x^{\star}\in\mathbb{R}^{n}$ to problem (1). Our aim is to iteratively find it via a distributed algorithm. Namely, given the iteration index $t\in\mathbb{N}$ and by denoting with $w_{i}^{t}\in{\mathbb{R}}^{n}$ the $i$-th agent’s estimate, at iteration $t$, of the solution to problem (1), our goal is to design a distributed protocol able to steer all these estimates to the minimizer $x^{\star}$. The peculiar challenge of this paper is that each agent $i$ can access neither the gradients $\nabla f_{i}(w_{i}^{t})$ (as in standard gradient-based methods) nor the cost functions in arbitrary points (as in standard zeroth-order methods). More in detail, we assume that agent $i$ can only use the single measurement $f_{i}(w_{i}^{t})$ properly combined with the so-called dither signal $d_{i}^{t}\in{\mathbb{R}}^{n}$ and the amplitude parameter $\delta>0$. The role of $d_{i}^{t}$ and $\delta$ will become clearer in the next section.
@@ III Extremum Seeking Tracking:
$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{r}}}\!+\!\phi_{i_{r}}\right)\!=\!0,$ (3c)
for $i\in\{1,\ldots,N\}$. Here, $\tau_{\text{per}}\in\mathbb{N}$ is the least common multiple of all periods $\tau_{ip}$. Sinusoidal dither functions are useful in practical applications to guarantee smooth inputs to the plant. However, other possibilities as, e.g., as square or triangular waves, are possibile [37].
@@
The main distinctive features of our method are as follows. First, errors in our method are given by third-order residuals as opposed to second-order ones in finite-difference methods. Second, gradient estimation is based on a single function query. This could be advantageous in scenarios in which multiple queries are expensive or even not allowed. Third and final, the estimation policy updates, and so the convergence guarantees, are purely deterministic. $\square$
@@
$\mathcal{D}_{i,r}\!:=\!\left\{\mathtt{col}(w_{i},z_{i})\in{\mathbb{R}}^{2n}\!\mid\!\left\|w_{i}-x^{\star}\right\|\leq r,s_{i}=2f_{i}(w_{i})/\delta\right\}\!.$
@@
Consider (4) and let Assumptions 12, and 3 hold. Then, for any $r,\bar{\rho}>0$, there exist $\gamma^{\star},\delta^{\star},k_{1}>0$, $\epsilon\in(0,\bar{\rho}/2)$, and $k_{2}\geq(\bar{\rho}/2-\epsilon)$ such that, for any $\gamma\in(0,\gamma^{\star})$ and $\delta\in(0,\delta^{\star})$, if $\mathtt{col}(w_{i}^{0},s_{i}^{0})\in\mathcal{D}_{i,r}$ for all $i\in\{1,\ldots,N\}$, the trajectories of (4) are bounded and for all $i\in\{1,\ldots,N\}$
@@
for all $t\geq t^{\star}$, where $t^{\star}\!:=\!-\frac{1}{\gamma k_{1}}\ln((\bar{\rho}/2\!-\!\epsilon)/k_{2})$, i.e., the convergence to the set $\{w_{i}\in{\mathbb{R}}^{n}\mid\left\|w_{i}^{t}\!-\!x^{\star}\right\|\leq\bar{\rho}\}$ is linear. $\square$
@@ IV Extremum Seeking Tracking:
where we decomposed $x$ according to $x:=\mathtt{col}(x_{1},\dots,x_{N})$ with $x_{i}\in{\mathbb{R}}^{n}$ for all $i\in\{1,\ldots,N\}$. We point out that, see also Fig. 1, system (9) can be conceived as an extremum-seeking scheme with output map $f(x+\delta d^{t})$.
@@
    We perform a first change of variables to describe the dynamics (9) in terms of the mean value (over the agents) $\tilde{z}_{\text{avg}}$ of $z$ and the orthogonal part $\tilde{z}_{\perp}$. Then, by relying on averaging theory (see [39] for discrete-time systems or [40, Ch.10] and [41] for continuous-time ones), we introduce a suitable auxiliary system named averaged system. The latter is obtained by averaging the original algorithm dynamics over a common period. The averaged system is shown to be driven by the cost gradients with additive estimation errors.
@@ IV-A Coordinate changes and averaged system
$\phi_{xz}(t,\tilde{x},\tilde{z})\!\!:=\!\!\begin{bmatrix}-L\tilde{x}-\tilde{z}-f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -L\tilde{z}-L(f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}\!,$
@@
where we have used the property $L\mathbf{1}=0$. As in [38], we take advantage of the initialization $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$. To this end, we introduce the novel coordinates $z_{\text{m}}\in{\mathbb{R}}^{n}$ and $\tilde{z}_{\perp}\in{\mathbb{R}}^{(N-1)n}$ representing the average of the variables $z_{i}^{t}$ (over the agents) and the orthogonal counterpart, namely
@@
The equation (16b) implies that $\tilde{z}_{\text{m}}^{t}=\tilde{z}_{\text{m}}^{0}$ for all $t\geq 0$. Then, since $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$ and $\mathbf{1}^{\top}G(\mathbf{1}x^{\star})=0$, it holds $\tilde{z}_{\text{m}}^{0}=0$ which allows us to ignore (16b) and rewrite (16) according to the equivalent, reduced system
@@
where $\phi:\mathbb{N}\times{\mathbb{R}}^{n_{\xi}}\to{\mathbb{R}}^{n_{\xi}}$ is introduced to compactly describe system (16a) with $\tilde{z}_{\text{m}}^{t}=0$ for all $t\geq 0$, i.e., $\phi$ is defined as
@@
For all $i\in\{1,\ldots,N\}$, there exists $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that, for any given $x_{i}\in{\mathbb{R}}^{n}$ and all $t\geq 0$, it holds
@@
for all $x\in{\mathbb{R}}^{Nn}$ and $t\geq 0$. Hence, by introducing $\xi_{\text{avg}}:=\mathtt{col}(\tilde{x}_{\text{avg}},\tilde{z}_{\perp,_{\text{avg}}})$ and by combining the definition of $\phi_{\text{avg}}$ in (21), $\phi$ in (19), and (25), we obtain
$\phi_{\text{avg}}(\xi_{\text{avg}})$ $=\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}-R^{\top}L(G(\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}$
@@
$\phi_{GT}(\xi_{\text{avg}})$ $:=\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}-R^{\top}L(G(\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}$

### v6

@@ Abstract
In this paper, we deal with a network of agents that want to cooperatively minimize the sum of local cost functions depending on a common decision variable. We consider the challenging scenario in which objective functions are unknown and agents have only access to local measurements of their local functions. We propose a novel distributed algorithm that combines a recent gradient tracking policy with an extremum seeking technique to estimate the global descent direction. The joint use of these two techniques results in a distributed optimization scheme that provides arbitrarily accurate solution estimates through the combination of Lyapunov and averaging analysis approaches with consensus theory. We perform numerical simulations in a personalized optimization framework to corroborate the theoretical results.
@@ I Introduction
We organize the literature review in two blocks: collaborative/distributed extremum seeking and distributed zeroth-order/derivative-free optimization. Early references on distributed extremum seeking for the so-called consensus optimization framework are [8] for a discrete-time setting and [9] for a continuous-time one. A proportional-integral extremum seeking design technique is proposed in [10], while, in [11], the gradient is approximated through a real-time protocol. In [12], authors propose the use of the sliding mode to generate the dither signal at the base of the extremum seeking. More recently, in [13], a distributed stochastic extremum seeking scheme is proposed for a source localization problem, while in [14] a distributed scheme approximating a Newton method is proposed. As for constraint-coupled distributed optimization, in [15], a distributed extremum seeking control, based on evolutionary game theory, is designed for real-time resource allocation. In [16], instead, a distributed, continuous-time scheme based on sign-based consensus is designed. In [17], a Lie bracket technique and extremum seeking are used for problems with linear constraints. In [18], resource allocation problems are addressed by an extremum seeking scheme.
@@
As for distributed zeroth-order/derivative-free optimization, authors in [19] develop a zeroth-order scheme based on a 1-point estimator and a gradient tracking policy. The work [20] instead proposes a zeroth-order algorithm based on a two-point estimator with a distributed gradient descent strategy and another one based on an $n$-point estimator with a gradient tracking policy. Authors in [21] propose a continuous-time gradient-free approach emulating a distributed gradient algorithm for which optimal asymptotic convergence is guaranteed. In [22], a sampled version of [21] is proposed. In [23], a continuous-time distributed algorithm based on random gradient-free oracles is proposed for convex optimization problems. In [24], the randomized gradient-free oracles introduced in [23] are used to build a gradient-free distributed algorithm in directed networks. Randomized gradient-free algorithms are used also in [25], where sequential Gaussian smoothing is used for non-smooth distributed convex constrained optimization. In [26], gradient-free optimization is addressed with the additional constraint that each agent can only transmit quantized information. The work in [27] instead develops a “directed-distributed projected pseudo-gradient” descent method for directed graphs. Paper [28] combines the gradient-free strategy of [23] with a saddle-point algorithm. Authors in [29] address an online constrained optimization problem by relying on the Kiefer-Wolfowitz algorithm to approximate the gradients, and [30] combines the estimation of the gradient via a “simultaneous perturbation stochastic approximation” technique with the so-called matrix exponential learning optimization method. In [31], a randomized gradient-free method is combined with a state-of-the-art distributed gradient descent approach for directed networks. In [32], an overview of zeroth-order methods based on a Frank-Wolfe framework is provided. Authors in [33] propose a distributed random gradient-free protocol to solve constrained optimization problems by using projection techniques. All the cited papers but [30] prove stability of the proposed schemes by using arguments not based on Lyapunov theory. We also note that [30] proposes an algorithm neither based on gradient tracking nor using extremum seeking.
@@
This paper proposes a novel algorithm to solve a distributed optimization problem in which network agents can only evaluate their local cost function at a given point, but not its gradient. The proposed solution consists of a distributed protocol in which the (unavailable) local gradients are approximated through an extremum seeking scheme. The approximations of the gradients are used to feed a suitable tracking mechanism which, in turn, allows for steering the local solution estimates along approximations of the global descent direction. The distributed algorithm uses a further consensus action on the local solution estimates. The convergence of our scheme is proved through Lyapunov and averaging theories for discrete-time systems. It is worth mentioning that our scheme, together with the ones in [34, 35], is the only distributed extremum seeking scheme proposed in discrete-time with the following distinctive features. The work [34] (i) does not address a consensus optimization problem and (ii) relies on consensus dynamics estimating the global cost, while ours estimates the global gradient. Instead, in [35], (i) the addressed consensus optimization problems have scalar decision variables and (ii) an extremum seeking technique is combined with a distributed gradient algorithm, i.e., without a tracking mechanism. As for the literature on distributed zeroth-order/derivative-free methods, the closest work is [19] in which extremum seeking is not used and the global gradient is approximated via a randomized 1-point policy. Consistently with the comparison table provided in [19], we highlight that there are no other distributed algorithms in the literature using 1-point gradient estimators. Indeed, although the distributed scheme in [36] uses 1-point gradient estimators, we remark that it is tailored for partition-based optimization, i.e., a simplified setup in which each local function depends only on the neighbors’ decision variables, see [3]. Also, our Lyapunov-based tools combined with averaging theory for discrete-time systems represent a distinctive feature in the algorithm analysis.
@@ II Problem Formulation
Each cost $f_{i}$ is $\mathcal{C}^{3}$ and has $\overline{L}_{i}$-Lipschitz continuous gradients. We denote $\overline{L}:=\max\{\overline{L}_{1},\dots,\overline{L}_{N}\}$. $\square$
Assumption 2 ensures the existence of a unique solution $x^{\star}\in\mathbb{R}^{n}$ to problem (1). Our aim is to iteratively find it via a distributed algorithm. Namely, given the iteration index $t\in\mathbb{N}$ and by denoting with $w_{i}^{t}\in{\mathbb{R}}^{n}$ the $i$-th agent’s estimate, at iteration $t$, of the solution to problem (1), our goal is to design a distributed protocol able to steer all these estimates to the minimizer $x^{\star}$. The peculiar challenge of this paper is that each agent $i$ cannot access either the gradients $\nabla f_{i}(w_{i}^{t})$ (as in standard gradient-based methods) or the cost functions in arbitrary points (as in standard zeroth-order methods). More in detail, we assume that agent $i$ can only use the single measurement $f_{i}(w_{i}^{t})$ properly combined with the so-called dither signal $d_{i}^{t}\in{\mathbb{R}}^{n}$ and the amplitude parameter $\delta>0$. The role of $d_{i}^{t}$ and $\delta$ will become clearer in the next section.
@@ III Extremum Seeking Tracking:
$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{q}}}\!+\!\phi_{i_{q}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{r}}}\!+\!\phi_{i_{r}}\right)\!=\!0,$ (3c)
for all $i\in\{1,\ldots,N\}$. Here, $\tau_{\text{per}}\in\mathbb{N}$ is the least common multiple of all periods $\tau_{i_{p}}$. Sinusoidal dither functions are useful in practical applications to guarantee smooth inputs to the plant. However, other possibilities, e.g., as square or triangular waves, are possibile [37].
@@
The main distinctive features of our method are as follows. First, errors in our method are given by third-order residuals as opposed to second-order ones in finite-difference methods. Second, gradient estimation is based on a single-function query per agent. This could be advantageous in scenarios in which multiple queries per agent are expensive or even not allowed. Third and final, the estimation policy updates, and so the convergence guarantees, are purely deterministic. $\square$
@@
$\mathcal{D}_{i,r}\!:=\!\left\{\mathtt{col}(w_{i},s_{i})\in{\mathbb{R}}^{2n}\!\mid\!\left\|w_{i}-x^{\star}\right\|\leq r,s_{i}=2f_{i}(w_{i})/\delta\right\}\!.$
@@
Consider (4) and let Assumptions 12, and 3 hold. Then, for any $r,\bar{\rho}>0$, there exist $\gamma^{\star},\delta^{\star},k_{1}>0$, $\epsilon\in(0,\bar{\rho}/2)$, and $k_{2}\geq(\bar{\rho}/2-\epsilon)$ such that, for any $\gamma\in(0,\gamma^{\star})$, $\delta\in(0,\delta^{\star})$, and $\mathtt{col}(w_{i}^{0},s_{i}^{0})\in\mathcal{D}_{i,r}$ for all $i\in\{1,\ldots,N\}$, the trajectories of (4) are bounded and satisfy
@@
for all $i\in\{1,\ldots,N\}$ and $t\geq t^{\star}:=\!-\frac{1}{\gamma k_{1}}\ln((\bar{\rho}/2\!-\!\epsilon)/k_{2})$, i.e., the convergence to the set $\{w_{i}\in{\mathbb{R}}^{n}\mid\left\|w_{i}^{t}\!-\!x^{\star}\right\|\leq\bar{\rho}\}$ is linear. $\square$
@@ IV Extremum Seeking Tracking:
where we decomposed $x$ according to $x:=\mathtt{col}(x_{1},\dots,x_{N})$ with $x_{i}\in{\mathbb{R}}^{n}$ for all $i\in\{1,\ldots,N\}$. We point out that system (9) can be conceived as an extremum seeking scheme with output map $f(x+\delta d^{t})$, see also Fig. 1.
@@
    We perform two change of variables to describe the dynamics (9) in terms of the mean value (over the agents) $\tilde{z}_{\text{avg}}$ of $z$ and the orthogonal part $\tilde{z}_{\perp}$ (both in error coordinates). Then, by relying on averaging theory (see [39] for discrete-time systems or [40, Ch.10] and [41] for continuous-time ones), we introduce a suitable auxiliary system named averaged system. The latter is obtained by averaging the original algorithm dynamics over a common period. The averaged system is shown to be driven by the cost gradients with additive estimation errors.
@@ IV-A Coordinate changes and averaged system
$\phi_{xz}(t,\tilde{x},\tilde{z})\!\!:=\!\!\begin{bmatrix}-L\tilde{x}-\tilde{z}-f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -L\tilde{z}-L(f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}\!.$
@@
where we have used the property $L\mathbf{1}=0$. As in [38], we take advantage of the initialization $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$. To this end, we introduce the novel coordinates $\tilde{z}_{\text{m}}\in{\mathbb{R}}^{n}$ and $\tilde{z}_{\perp}\in{\mathbb{R}}^{(N-1)n}$ representing the average of the variables $\tilde{z}_{i}^{t}$ (over the agents) and the orthogonal counterpart, namely
@@
The equation (16b) implies that $\tilde{z}_{\text{m}}^{t}=\tilde{z}_{\text{m}}^{0}$ for all $t\in\mathbb{N}$. Then, since $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$ and $\mathbf{1}^{\top}G(\mathbf{1}x^{\star})=0$, it holds $\tilde{z}_{\text{m}}^{0}=0$ which allows us to ignore (16b) and rewrite (16) according to the equivalent, reduced system
@@
where $\phi:\mathbb{N}\times{\mathbb{R}}^{n_{\xi}}\to{\mathbb{R}}^{n_{\xi}}$ is introduced to compactly describe system (16a) with $\tilde{z}_{\text{m}}^{t}=0$ for all $t\in\mathbb{N}$, i.e., $\phi$ is defined as
@@
For all $i\in\{1,\ldots,N\}$, there exists $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that, for any given $x_{i}\in{\mathbb{R}}^{n}$ and all $t\in\mathbb{N}$, it holds
@@
for all $x\in{\mathbb{R}}^{Nn}$ and $t\in\mathbb{N}$. Hence, by introducing $\xi_{\text{avg}}:=\mathtt{col}(\tilde{x}_{\text{avg}},\tilde{z}_{\perp,_{\text{avg}}})$ and by combining the definition of $\phi_{\text{avg}}$ in (21), $\phi$ in (19), and (25), we obtain
$\phi_{\text{avg}}(\xi_{\text{avg}})\!$ $=\!\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}\!-\!R^{\top}L(G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})\!-\!G(\mathbf{1}x^{\star}))\end{bmatrix}$
@@
$\phi_{GT}(\xi_{\text{avg}})\!$ $:=\!\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ \!-\!R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}\!-\!R^{\top}L(G(\tilde{x}_{\text{avg}}\!+\!\mathbf{1}x^{\star})\!-\!G(\mathbf{1}x^{\star}))\end{bmatrix}$